Find the standard equation with focus: (-2,0) and Directrix: x=4
(y-k)2 = -4p(x-h) This is as far as I have gotten. I think that the vertex is (1,0)

Answer by lwsshak3(11628)

Find the standard equation with focus: (-2,0) and Directrix: x=4
..
Since the focus is to the left of the directrix, this is a parabola which opens leftward with axis of symmetry being the x-axis or y=0. Standard form for this parabola is:
y^2=-4px
Distance between focus on the axis of symmetry=6
p=1/2 this distance = 3
vertex on the axis of symmetry is at (1,0) (located halfway between focus and directrix on axis of symmetry)
Equation of given parabola:
y^2=-12(x-1)
see the graph below for a visual representation of the equation for given parabola.
y=(-12(x-1))^.5